OpenStudy (neer2890):

for n belongs to N(Natural numbers),find the value of the integral:

OpenStudy (usman1995):

where's integral

OpenStudy (neer2890):

\[\int\limits_{0}^{n} \frac{ 1-(\frac{ x }{ n })^{n} }{ n-x } dx\]

OpenStudy (neer2890):

???? @Usman1995

OpenStudy (usman1995):

sorry i havnt studied integration yet

OpenStudy (neer2890):

ok. thanks though

OpenStudy (perl):

well for x = 1 , that is easy

OpenStudy (perl):